# AV-Reasoner: Improving and Benchmarking Clue-Grounded Audio-Visual Counting for MLLMs

### v1

@@ 5.2 Comparison with Base Model on Counting Tasks
Ola-Omni (Base Model) 16.50 51.50 36.10 30.91 44.44 17.92 38.85 4.57 10.52 16.94 37.10 4.54 10.59 0.42 75.66
AV-Reasoner-Thinking (Ours) 44.00 81.00 36.59 34.55 46.67 21.03 48.78 3.26 8.20 22.78 48.59 3.13 6.69 1.22 79.65
(+27.50) (+29.50) (+0.49) (+3.64) (+2.23) (+3.11) (+9.93) (-1.31) (-2.32) (+5.84) (+11.49) (-1.41) (-3.90) (+0.80) (+3.99)
AV-Reasoner (Ours) 43.50 80.50 37.07 32.12 47.78 22.30 48.30 3.15 5.89 23.56 48.78 3.14 6.12 0.92 80.82
(+27.00) (+29.00) (+0.97) (+1.21) (+3.34) (+4.38) (+9.45) (-1.42) (-4.63) (+6.62) (+11.68) (-1.40) (-4.47) (+0.50) (+5.16)
@@ 5.3 Ablation Studies
Base model 82.61 79.47 19.43 14.5 14.5 0.28 0.19 16.50 17.92 16.94 0.42 36.10 30.91 44.44
CB-RL(w/ SRM) + FT-RL 92.97 84.59 78.56 51.90 49.80 55.45 45.55 29.50 18.31 18.99 0.84 36.59 31.71 45.37
SFT 78.73 78.10 80.02 48.70 47.10 52.67 45.48 41.50 15.00 15.97 0.60 27.80 30.91 35.52
SFT + GRPO with all data 93.25 85.68 79.82 58.10 55.80 60.09 48.43 31.50 10.42 10.91 0.55 24.39 22.42 32.20
SFT + CB-RL(w/o SRM) 89.24 78.93 80.03 60.30 58.30 63.91 45.97 42.50 20.84 21.03 0.86 34.63 30.73 42.44
SFT + CB-RL(w/o SRM) + FT-RL 93.13 83.44 80.10 61.30 60.90 64.28 46.72 43.50 20.74 21.42 1.16 35.61 33.17 46.34
SFT + CB-RL(w/ SRM) 93.14 84.79 80.15 64.50 61.60 65.07 46.62 43.00 20.84 21.62 1.20 35.12 33.66 45.85
SFT + CB-RL(w/ SRM) + FT-RL 93.17 85.01 81.26 64.70 62.00 65.18 46.73 44.00 21.03 22.78 1.22 36.59 34.55 46.67

### v2

@@ 5.2 Comparison with Base Model on Counting Tasks
Ola-Omni (Base Model) 16.50 51.50 36.10 30.91 44.44 17.92 38.85 4.57 10.52 25.33 46.53 3.30 5.98 0.84 75.66
AV-Reasoner-Thinking (Ours) 44.00 81.00 36.59 34.55 46.67 21.03 48.78 3.26 8.20 34.08 60.95 2.40 4.66 1.68 79.65
(+27.50) (+29.50) (+0.49) (+3.64) (+2.23) (+3.11) (+9.93) (-1.31) (-2.32) (+8.75) (+14.42) (-0.90) (-1.32) (+0.84) (+3.99)
AV-Reasoner (Ours) 43.50 80.50 37.07 32.12 47.78 22.30 48.30 3.15 5.89 35.83 61.44 2.38 4.44 1.11 80.82
(+27.00) (+29.00) (+0.97) (+1.21) (+3.34) (+4.38) (+9.45) (-1.42) (-4.63) (+10.50) (+14.91) (-0.92) (-1.54) (+0.27) (+5.16)
@@ 5.3 Ablation Studies
Base model 82.61 79.47 19.43 14.5 14.5 0.28 0.19 16.50 17.92 25.33 0.84 36.10 30.91 44.44
CB-RL(w/ SRM) + FT-RL 92.97 84.59 78.56 51.90 49.80 55.45 45.55 29.50 18.31 28.41 1.15 36.59 31.71 45.37
SFT 78.73 78.10 80.02 48.70 47.10 52.67 45.48 41.50 15.00 23.89 0.83 27.80 30.91 35.52
SFT + GRPO with all data 93.25 85.68 79.82 58.10 55.80 60.09 48.43 31.50 10.42 17.16 0.77 24.39 22.42 32.20
SFT + CB-RL(w/o SRM) 89.24 78.93 80.03 60.30 58.30 63.91 45.97 42.50 20.84 33.07 1.19 34.63 30.73 42.44
SFT + CB-RL(w/o SRM) + FT-RL 93.13 83.44 80.10 61.30 60.90 64.28 46.72 43.50 20.74 33.68 1.61 35.61 33.17 46.34
SFT + CB-RL(w/ SRM) 93.14 84.79 80.15 64.50 61.60 65.07 46.62 43.00 20.84 34.00 1.66 35.12 33.66 45.85
SFT + CB-RL(w/ SRM) + FT-RL 93.17 85.01 81.26 64.70 62.00 65.18 46.73 44.00 21.03 34.08 1.68 36.59 34.55 46.67